OpenStudy (anonymous):

The length of the hypotenuse of a 30°-60°-90° triangle is 12. Find the perimeter.

OpenStudy (anonymous):

It might be about \[12+6+6\sqrt{3}=18+6\sqrt{3}\]

OpenStudy (anonymous):

I thought that's what it would be, but I just wanted to double check and see if someone else got the same answer. thank you

OpenStudy (anonymous):

In trigspeak, we could say it is\[P=12(1+\cos \theta+\sin \theta)\]\[\theta=\pi/6\]